OpenStudy (abdela25):

quick help with a math word problem MEDALS WILL BE REWARDED read full replys before answering

OpenStudy (anonymous):

what is the question?

OpenStudy (abdela25):

This is the word problem the debate club needs $240.00 to attend a debate tournament the club decides to sell cups of iced tea and lemonade at baseball games. iced tea will be sold for $0.50 and lemonade will be sold for $0.80 per cup

OpenStudy (anonymous):

just add and you will gt the amount of $240.00

OpenStudy (abdela25):

the first question to it is write an equation to find how many cups of each bevergae must be sold to rais $240.00 My answer to this one was: x=number of cups of iced tea y=number of cups of lemonade 0.5x+0.8y=240, 5x+8y+2400, 8y+-5x+2400

OpenStudy (anonymous):

yes that is correct

OpenStudy (abdela25):

the second question to this problem was: Graph the equation. what are the x- and y-intercepts my answer was: x-intercept- y=0, -5/8x+300=0, -5/8x=-300 x=-300/(-5/8), x=-300(-8/5), x=2400/5 x=480, x-intercept = (480,0) y-intercept- x=0, y=-5/8(0)+300, y=300 y-intercept = (0,300)

OpenStudy (abdela25):

but this part of the question also acts for a graph i dont know how to graph this

OpenStudy (anonymous):

idk either

OpenStudy (anonymous):

um try google

OpenStudy (abdela25):

google isnt helpful with this graphing stuff

OpenStudy (anonymous):

yea ig ill try to figure it out

OpenStudy (abdela25):

ok thanks

OpenStudy (anonymous):

yw